OpenStudy (anonymous):

4sqrt{80} 4sqrt{80} @Mathematics

OpenStudy (anonymous):

\[\sqrt[4]{}80\]

OpenStudy (anonymous):

\[\sqrt[4]{80} = 2\sqrt[4]{5}\]

OpenStudy (anonymous):

would the factorization be 8x10?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

16*5

OpenStudy (anonymous):

first tena u learn how to use equation editor

OpenStudy (anonymous):

i am i took out the ?

OpenStudy (anonymous):

yeah u have in the second case only

OpenStudy (anonymous):

im learning as i go, first time ive used it